# A Survey of LLM-Driven AI Agent Communication: Protocols, Security Risks, and Defense Countermeasures

### v3

@@ IX-A1 Powerful but Lightweight Malicious Input Filter
We deem that user inputs are still the largest-scale attack carrier in the agent ecosystem, especially considering that the inputs are becoming more open (no longer limited to user instructions but also contain environment feedback), multimodal, and semantically complex. Besides, the future agent ecosystem will pay more attention to effectiveness, especially given that the running speed of LLMs is inherently slow. Such dual demand will put a very heavy burden on related defenses. As a result, to mitigate this problem, lightweight but powerful malicious input filters must be established. This not only requires mature techniques in AI to slim the defense models down (just like DeepSeek), but also needs to integrate with other techniques, such as offloading some fundamental computing on the programmable line-speed devices (e.g., programmable switches and SmartNICs) to facilitate the input filtering process.
@@ IX-B2 Protect Intellectual Property Rights
Nowadays, there has been a large number of LLMs that have been open-sourced. These can act as the brain of different agents. However, even for open-source LLMs, the publishers still restrict their application scope, e.g., other developers should also open-source their agents built on these LLMs. However, there still lacks laws to effectively protect such intellectual property. For example, the criteria for determining plagiarism in agents is not clear. Even if plagiarism is determined, there is still a lack of defining standards for the degree of plagiarism (e.g., 50% or 90%?). We think there urgently need related laws and regulations.
@@ IX-B3 Cross-border Supervision
To our knowledge, the related formulation of laws and regulations (such as those related to agent crimes) lags far behind the development of agents. For example, how to define the theft and misappropriation of agents, the accident responsibility of autonomous driving agents
@@ X Conclusion
This survey systematically reviews the security issues of agent communication. We first highlight the differences between previous related surveys and this survey, and summarize the preliminaries of LLM-driven agents. Then, we make a definition and classification of agent communication to help future researchers quickly classify and evaluate their work. Next, we detailedly illustrate the communication protocols, security risks, and possible defense countermeasures for three agent communication stages, respectively. Then, we conduct experiments using MCP and A2A to help illustrate the new attack surfaces brought by agent communication. Finally, we discuss the open issues and future directions from technical and legal aspects, respectively.

### v4

@@ IX-A1 Powerful but Lightweight Malicious Input Filter
We deem that user inputs are still the largest-scale attack carrier in the agent ecosystem, especially considering that the inputs are becoming more open (no longer limited to user instructions but also containing environment feedback), multimodal, and semantically complex. Besides, the future agent ecosystem will pay more attention to effectiveness, especially given that the running speed of LLMs is inherently slow. Such dual demand will put a very heavy burden on related defenses. As a result, to mitigate this problem, lightweight but powerful malicious input filters must be established. This not only requires mature techniques in AI to slim the defense models down (just like DeepSeek), but also needs to integrate with other techniques, such as offloading some fundamental computing on the programmable line-speed devices (e.g., programmable switches and SmartNICs) to facilitate the input filtering process.
@@ IX-B2 Protect Intellectual Property Rights
Nowadays, there has been a large number of LLMs that have been open-sourced. These can act as the brain of different agents. However, even for open-source LLMs, the publishers still restrict their application scope, e.g., other developers should also open-source their agents built on these LLMs. However, there still lack laws to effectively protect such intellectual property. For example, the criteria for determining plagiarism in agents is not clear. Even if plagiarism is determined, there is still a lack of defining standards for the degree of plagiarism (e.g., 50% or 90%?). We think there urgently need related laws and regulations.
@@ IX-B3 Cross-border Supervision
To our knowledge, the related formulation of laws and regulations (such as those related to agent crimes) lags far behind the development of agents. For example, how to define the theft and misappropriation of agents, and the accident responsibility of autonomous driving agents.
@@ X Conclusion
This survey systematically reviews the security issues of agent communication. We first highlight the differences between previous related surveys and this survey, and summarize the evolution direction of communication technology. Then, we make a definition and classification of agent communication to help future researchers quickly classify and evaluate their work. Next, we detailedly illustrate the communication protocols, security risks, and possible defense countermeasures for three agent communication stages, respectively. Then, we conduct experiments using MCP and A2A to help illustrate the new attack surfaces brought by agent communication. Finally, we discuss the open issues and future directions from technical and legal aspects, respectively.